As of November 2, 2026, your Los Alamos apiary faces an urgent stores situation: Iris is showing the relapse signature — her mean climbed to 84.8°F (29.3°C) two days ago then fell back to 74.3°F (23.5°C), the classic pattern of a dose burned through and cells bare again — and she needs bulk feeding today, not a dose. Hive #2 and Hive #1 are holding solid brood temperatures with strong acoustics, and Hive #3 remains thermally stable with tight damping, but the apiary-wide advisory is firm: 6 of the last 7 days have offered little or no foraging weather, temperature and vibration cannot tell you what is capped in any cell, and every hive needs a heft this week.
The relapse arc on Iris — mean up sharply, held briefly, now falling again with damping retained — matches the El Viento 2026-10 pattern exactly: a populous colony keeps its amplitude tight while running down fuel, so the trajectory of the mean (not the swing) is the diagnostic signal (Minaud et al. 2024, Ecol. Indic.; Colin et al. 2025, Ecol. Inform.).
Hive #1, Hive #2, and Hive #3 each maintain tight 1.6–2.6°F (0.9–1.4°C) diurnal swings anchored ~84–86°F (28–30°C) above the 45.1°F (7.3°C) outdoor reference, confirming brood-core damping fully intact; Iris shows a 3.2°F (1.8°C) swing at a 74.3°F (23.5°C) mean — damping is present but the mean has now fallen 12.3°F (6.9°C) below her all-time average, a falling trajectory with retained tight amplitude that is the relapse signature.
-
Iris critical Score: 38/100📋 Overall Health
Your Iris hive is showing the relapse signature: her mean climbed to 84.8°F (29.3°C) two days ago, held briefly, and has now fallen back to 74.3°F (23.5°C) — a −10.5°F (−5.9°C) drop while the daily swing remains tight at 3.2°F (1.8°C) and a warm offset above the 45.1°F (7.3°C) outdoor reference is still present; this is a populous colony burning through its last fuel, not a brood break, and it needs bulk feeding today. The sustained 17-day acoustic step-down (all bands 5–8 dB below her own baseline) continues unchanged, consistent with a reduced brood-rearing footprint; humidity is healthy at 56.4%; vibration kurtosis is 2.7 (normal, no impact events); watch whether the mean stabilizes or continues to slide after feeding as your key forward indicator.
🌅 Diurnal Cycle PatternsIris's 24h temperature arc is flat and narrow — 72.7–75.9°F (22.6–24.4°C), swing 3.2°F (1.8°C) — with a slow falling trend over the latest 12 hours (−2.0°F / −1.1°C); the tight amplitude with a falling mean is the relapse fingerprint (retained damping, depleting fuel). Gas resistance at 481 kΩ is above her 384 kΩ all-time average, and the heater-scan curve shows a shape broadly similar to previous readings with a minor low-temperature rhythm shift (dominant daily cycle migrating from 150°C to 100°C steps) consistent with changing humidity or propolis rather than a new chemistry event; vibration shows a weak daily rhythm only in the 5–20 Hz band, and no impact events.
🔬 Correlations & Discoveriestemperature vs gas_resistance (14d) | r = -0.93The partial correlation of gas resistance with in-hive temperature (controlling for weather, r=−0.93) is the strongest signal in Iris's dataset: as her brood footprint contracts and the cluster cools, brood pheromone load (E-β-ocimene and related volatiles) falls, reducing the MOX sensor's reducing-gas exposure and allowing resistance to rise — confirming the temperature decline reflects a genuine reduction in brood-rearing activity rather than a sensor artifact.📚 Reference: Stabentheiner et al. 2010, PLoS ONE, Maisonnasse et al. 2010, PLoS ONE, Seeley 1985temperature vs sound_spectrum (14d) | r = -0.85The sustained 17-day parallel decline in both mean temperature and acoustic energy (all bands 5–8 dB below Iris's own baseline) with retained tight daily amplitude is the relapse pattern documented at El Viento 2026-10: a colony still producing heat but running low on fuel shows a falling mean with intact damping, not an expanding swing — the trajectory of the mean is the diagnostic, and it continues to fall.📚 Reference: Minaud et al. 2024, Ecol. Indic., Colin et al. 2025, Ecol. Inform., Fahrenholz et al. 1989, J. Comp. Physiol. B⚠️ Issue: RELAPSE SIGNATURE — stores critically low: mean rose to 84.8°F (29.3°C) two days ago then fell back to 74.3°F (23.5°C), a −10.5°F (−5.9°C) drop with damping retained — a dose has been consumed and cells are likely bare again; bulk feeding needed today⚠️ Issue: Sustained 17-day acoustic step-down: all 10 spectral bands 5–8 dB below Iris's own all-time baseline, consistent with a contracted or broodless nest▹ Action: Feed Iris by WEIGHT today — bulk 2:1 syrup (2 kg sugar : 1 L water) placed directly over the cluster while daytime highs stay above 50°F (10°C); switch to fondant or a sugar brick on the top bars when it stays colder; a single dose will produce another brief warm-up arc — keep feeding until the hive hefts heavy toward a 40–50 lb gross stores target, never stop at one dose▹ Action: After feeding, monitor whether the mean temperature stabilizes and holds above 80°F (27°C) for more than 48 hours — a second relapse (mean rising then falling again) means the cluster is still under-fuelled; a sustained recovery means stores are rebuilding▹ Action: On your next visit also assess queen presence and brood pattern — the 17-day acoustic step-down and cooling drift are consistent with a brood-break or requeening gap, and confirming queen status remains important, but stores come first -
Hive #2 warning Score: 72/100📋 Overall Health
Your Hive #2 averaged 92.1°F (33.4°C) today — essentially matching her all-time average of 92.4°F (33.5°C) — with a tight 2.6°F (1.4°C) daily swing and a robust ~47°F (~26°C) warm offset above the 45.1°F (7.3°C) outdoor reference, confirming solid active thermoregulation just inside the ideal 93–97°F (34–36°C) band; the 14-day acoustic rising trend continues (+4.7 dBFS), with dominant energy now in the 200–260 Hz colony-fundamental band at −25.2 dB (4.7–5.3 dB above her own baseline in the low bands), a thermally and acoustically vigorous single-deep heading into November. Humidity averaged 46.8%, consistent with dry Los Alamos November air through the screened floor; vibration kurtosis is 2.9 (normal); battery is draining moderately (−64 mV today, current 3,976 mV, still well above the 3,796 mV all-time low); heft this week is the priority — single-deep configuration means the smallest reserve buffer in the apiary.
🌅 Diurnal Cycle PatternsHive #2's 24h arc is 91.4–94.0°F (33.0–34.5°C), swing 2.6°F (1.5°C), stable trend — textbook brood-core damping. The acoustic spectrum shows 9/10 bands with a clear daily cycle (strongest in 100–150 Hz at 3.7 dB rhythm) and a notable spectral evolution: dominant daily-rhythm band has shifted from 1500–3600 Hz to 150–200 Hz over the window, and band energy in 100–260 Hz has risen +4.7–5.3 dB relative to baseline, consistent with increased low-frequency colony hum from a growing or highly active cluster. Gas heater-scan shows a shifted shape at low plate temperatures (100°C step +37%, 150°C +33%, 200°C +28% above prior levels) co-varying with the rising colony activity — likely propolis work, brood pheromone load, or processing activity, not a disease signal; vibration shows virtually no daily rhythm (max 1.0 dB in 5–20 Hz), with dominant energy in the 80–120 Hz structural resonance band.
🔬 Correlations & Discoveriestemperature vs humidity (14d) | r = +0.76The genuine deconfounded temperature–humidity coupling (r=+0.76, partial r=+0.57 controlling for weather) in Hive #2 reflects active hygroregulation: as the cluster expands metabolic output to heat the brood, respiratory water vapor and fanning effort co-vary, raising both temperature and local humidity together — the textbook signature of a brood-rearing colony actively managing its microclimate.📚 Reference: Human et al. 2006, Naturwissenschaften, Ellis et al. 2010, Naturwissenschaften, Eouzan et al. 2019, PLOS ONEsound_spectrum vs gas_resistance (14d) | r = -0.65The rising low-frequency acoustic energy (+4.7–5.3 dB in the 100–260 Hz colony-fundamental bands) alongside falling gas resistance at low heater-scan steps (100–200°C) is consistent with an increasingly active colony producing more brood pheromones, propolis volatiles, and fanning-circulated terpenes — all benign correlates of a colony ramping up rather than a stress event.📚 Reference: Maisonnasse et al. 2010, PLoS ONE, Bankova et al. 2000, Apidologie, Ramsey et al. 2020, Sci. Rep.⚠️ Issue: Stores unconfirmed: single-deep configuration carries the smallest reserve buffer in the apiary; temperature and vibration cannot confirm what is capped in the cells, and 6 of the last 7 days offered little or no foraging weather▹ Action: Heft Hive #2 this week — her single-deep capacity means the smallest margin before stores run out; if light, feed by weight toward at least 40–50 lb gross using 2:1 syrup placed over the cluster while daytime highs stay above 50°F (10°C), then switch to fondant or a sugar brick on the top bars when it stays colder -
Hive #1 healthy Score: 88/100📋 Overall Health
Your Hive #1 averaged 92.1°F (33.4°C) today — 1.2°F (0.6°C) below her all-time average of 93.3°F (34.0°C) but solidly inside the ideal 93–97°F (34–36°C) band — with a remarkably tight 1.8°F (1.0°C) daily swing, a ~47°F (~26°C) warm offset above the 45.1°F (7.3°C) outdoor reference, and a sound spectrum classified near-stable (all bands within 0.5–2.7 dB of her own baseline, max deviation only 2.7 dB), making this the most thermally and acoustically consistent colony in the apiary heading into November. Humidity averaged 46.1%, consistent with dry Los Alamos air; vibration kurtosis is 2.7 (normal); a mild 14-day cooling drift (−2.0°F / −1.1°C) and gentle acoustic softening (−1.6 dBFS over 14 days) with retained tight damping are consistent with normal autumn brood contraction, not a stress event; heft this week remains the single priority action.
🌅 Diurnal Cycle PatternsHive #1's 24h arc is 91.1–92.9°F (32.9–33.9°C), swing 1.8°F (1.0°C) — the tightest in the apiary and a strong indicator of active brood-core clamping. Sound shows 9/10 bands with a daily cycle (strongest in 1500–3600 Hz at 12.8 dB rhythm), with a modest spectral softening (260–350 Hz −2.4 dB, 150–200 Hz −2.3 dB over 14 days) consistent with gradual brood contraction; the dominant vibration daily rhythm has shifted from 5–20 Hz to 120–160 Hz over the window, suggesting fanning/ventilation activity is now more structured than earlier in the season. Gas heater-scan shows rising low-temperature step resistance (100°C +22%, 150°C +23%, 200°C +21%) alongside stable mid- and high-temperature steps, consistent with a modest increase in heavier volatile load (propolis, brood pheromones) rather than a chemistry change — the PCA colony-state index rising over the window corroborates overall stability.
🔬 Correlations & Discoveriestemperature vs gas_resistance vs humidity (14d) | r = +0.61The genuine deconfounded positive coupling of temperature with both gas resistance (r=+0.61) and humidity (r=+0.78) in Hive #1 — unusual in that gas resistance rises with temperature rather than falls — suggests that as the colony heats its brood core, it also increases ventilation efficiency (drying the local air and clearing VOCs), with the screened floor pulling dry November air through and reducing reducing-gas accumulation when colony metabolic activity is highest: a healthy, well-ventilated winter-preparation colony signature.📚 Reference: Kleinhenz et al. 2003, J. Exp. Biol., Stabentheiner et al. 2010, PLoS ONE, Maisonnasse et al. 2010, PLoS ONEsound_energy vs temperature (14d) | r = +0.45The partial sound–temperature correlation (r=+0.45 controlling for weather) confirms that Hive #1's acoustic energy tracks her own brood-nest activity rather than outdoor conditions — a tight, hive-specific coupling consistent with a well-regulated colony where sound energy rises and falls with the thermal footprint of active brood rearing, the expected autumn pattern as both contract gradually together.📚 Reference: Ramsey et al. 2020, Sci. Rep., Meikle et al. 2017, Apidologie, Minaud et al. 2024, Ecol. Indic.▹ Action: Heft Hive #1 this week as part of the apiary-wide November stores check — her excellent thermal and acoustic profile cannot confirm what is capped in the cells, and temperature cannot see stores; if light, feed by weight toward at least 40–50 lb gross using 2:1 syrup while daytime highs stay above 50°F (10°C), then switch to fondant or a sugar brick on the top bars when it stays colder -
Hive #3 warning Score: 65/100📋 Overall Health
Your Hive #3 averaged 93.4°F (34.1°C) today — 1.2°F (0.7°C) below her all-time average of 94.6°F (34.8°C) but inside the ideal 93–97°F (34–36°C) band — with a tight 1.3°F (0.7°C) daily swing and a ~48°F (~27°C) warm offset above the 45.1°F (7.3°C) outdoor reference, confirming brood-core damping fully intact; however, the 14-day acoustic softening has deepened to −7.3 dBFS overall with low bands (100–260 Hz) now 3.6–6.1 dB below her own baseline and all 9/10 bands trending down in the second half of today, a multi-week pattern that — while consistent with seasonal brood contraction — warrants a heft and a brief inspection to confirm brood coverage and queen status this week. Humidity averaged 42.7% (range 40.6–44.7%), persistently just above the 40% lower boundary driven by dry November outdoor air through the screened floor; vibration kurtosis is 3.3 (normal, near the 3.0 baseline) with the highest vibration RMS in the apiary (0.0206 g avg, strong daily peaks to 33.6 mG).
🌅 Diurnal Cycle PatternsHive #3's 24h arc is 92.7–94.0°F (33.7–34.4°C), swing 1.3°F (0.7°C) — tight brood damping fully intact. Acoustic spectrum shows a broad softening in all low and mid bands (100–420 Hz down 3–6 dB from all-time baseline) with all bands trending further down in the second half of today (100–800 Hz each falling 1.6–2.4 dB); the dominant daily rhythm lives in 1500–3600 Hz (15.2 dB amplitude), and a 14-day spectral shift away from lower frequencies (100–260 Hz −7.1 dB) suggests the colony's low-frequency hum footprint is contracting — consistent with brood reduction. Gas heater-scan shows low-temperature steps rising (100°C +19%, 150°C +16%, 200°C +14%) with a strengthening daily rhythm at the 150°C step (24% → 41% swing), likely reflecting propolis or metabolic water cycling; vibration shows strongest daily rhythm in 50–80 Hz (colony metabolic baseline band), and the dominant vibration band at 80–120 Hz is unchanged from baseline.
🔬 Correlations & Discoveriessound_energy vs temperature (14d) | r = +0.68The strongest partial correlation in Hive #3 is sound energy vs. in-hive temperature controlling for weather (r=+0.68), meaning acoustic energy tracks the thermal footprint of active brood-rearing independent of outdoor conditions — the shared 14-day cooling drift (−2.5°F / −1.4°C) and acoustic softening (−7.3 dBFS) are a tightly coupled autumn brood-contraction signal, and the change-point in temperature around October 22 anchors when this transition began.📚 Reference: Ramsey et al. 2020, Sci. Rep., Meikle et al. 2017, Apidologie, Minaud et al. 2024, Ecol. Indic.humidity vs gas_resistance (14d) | r = -0.62The genuine deconfounded humidity–gas resistance anti-correlation (r=−0.62) in Hive #3 is consistent with MOX sensor humidity confounding: as screened-floor ventilation pulls in drier November outdoor air, the sensor's MOX plate sees less water vapor and resistance rises slightly — the gas channel here reflects ventilation dynamics and the dry Los Alamos environment rather than a chemistry event, and the daily 100°C rhythm shift corroborates this moisture-driven mechanism.📚 Reference: Burgues & Marco 2018, Sensors, Abidin et al. 2022, Sensors, Seeley 1974, J. Insect Physiol.⚠️ Issue: Continued 14-day acoustic softening now at −7.3 dBFS overall (low bands 100–260 Hz down 3.6–6.1 dB from all-time baseline), with all bands trending further down in the second half of today alongside a −2.5°F (−1.4°C) temperature drift — the two signals remain coupled and warrant a brief inspection to confirm brood coverage and queen status⚠️ Issue: In-hive humidity averaging 42.7% (range 40.6–44.7%), persistently just above the 40% lower boundary — driven by dry outdoor air through the screened floor; passive monitoring note at current brood temperatures with the core holding 93–94°F (34–34.4°C)⚠️ Issue: Stores unconfirmed: highest activity and vibration level in the apiary means the fastest potential burn rate if stores are short; 6 of the last 7 days offered little or no foraging weather▹ Action: Heft Hive #3 this week — her highest activity level means the fastest potential burn rate if stores are short; if light, feed by weight toward at least 40–50 lb gross using 2:1 syrup while daytime highs stay above 50°F (10°C), then switch to fondant or a sugar brick on the top bars when it stays colder▹ Action: On this same visit, take a brief look at brood coverage and confirm queen status — the 14-day coupled acoustic and thermal softening is consistent with normal seasonal contraction, but after two more weeks of decline a visual check is the right call; damping remains fully intact so there is no urgency beyond folding this into the stores visit you are already making